Computer equipment was acquired at the beginning of the year at a cost of $63,500. It had an estimated residual value of $3,500 and an estimated useful life of five years. Determine the (a) depreciable cost, (b) straight-line rate, and (c) annual straight-line depreciation.

a. Depreciable cost $
b. Straight-line rate %
c. Annual straight-line depreciation $
Answer #1

Answers

  • Working

A

Cost

$            63,500.00

B

Residual Value

$              3,500.00

C=A - B

Depreciable base

$            60,000.00

D

Life [in years]

5

E=C/D

Annual SLM depreciation

$            12,000.00

F=E/C

SLM Rate

20.00%

  • Requirements

[a] Depreciable cost = $ 60000
[b] Straight Line rate = 20%
[c] Annual Straight Line Depreciation = $ 12,000